Consider an open economy that is described by the following model:

C = 200 + 0.75Y

Where:

Y = income

C = consumption

 

1.Calculate the equilibrium income.  

2. Calculate the amount of tax that the government collects at equilibrium.  

3.Calculate the government budget, and explain whether there is a surplus or a deficit.

4. Calculate the change in investment required to reach a full employment level of income.


1
Expert's answer
2021-03-24T20:33:10-0400

Consider an open economy that is described by the following model:

C = 200 + 0.75Y

Where:

Y = income

C = consumption

 

1.   Calculate the equilibrium income. 

The formula for the equilibrium level of income is when aggregate supply (AS) is equal to aggregate demand (AD), where AS = AD.

AS = AD

Y= 200 + 0.75Y + 200

Y – 0.75 Y = 200 + 200

0.25 Y = 400

Y = 1,600

2.     Calculate the amount of tax that the government collects at equilibrium. 

The tax revenue is calculated by multiplying the tax per unit by the total quantity sold Qt. The tax incidence on the consumers is given by the difference between the price paid Pc and the initial equilibrium price Pe,

1,600 – 200

The answer is 1,400

3.     Calculate the government budget, and explain whether there is a surplus or a deficit.

The government budget balance is the difference between revenues and expenditures. In this case, there will be a surplus.

1,600 – 400

The answer is 1,200

4. Calculate the change in investment required to reach a full employment level of income.

Take the selling price and subtract the initial purchase price. The result is the gain or loss. Take the gain or loss from the investment and divide it by the original amount or purchase price of the investment. Finally, multiply the result by 100 to arrive at the percentage change in the investment

1,600 – 400 = 1,200

1,200 ÷ 200 = 6

6 × 100 =600%